The rate of electron emission from 4mg of Pb (atomic numer=80, atomic mass=210) wwith half life of 5 days is?
1
Expert's answer
2020-06-11T10:35:48-0400

moles of Pb here "\frac{4\times10^{-3}}{210}=1.9\times10^{-5}"

moles of electron here = "80\times1.9\times10^{-5}" ="1.52\times10^{-3}"

"k=\frac{\ln2}{t_{1/2}}=\frac{0.69}{5}=1.38day^{-1}"

rate of electron emission = "kA=1.38\times1.52\times10^{-3}=2.1\times10^{-3}eletrons/day"